the sum of two numbers is 25. If 2/3 of the smaller one is 2 more than 1/4 of the larger find the two numbers
Answer by mananth(16949)

let one number be x
then other number = 25-x
2x/3 = 2+1/4*(25-x)
multiply by 12
8x= 24+3(25-x)
8x=24+75-3x
11x=99
x=9
9 & 16 are the numbers